The distance between 2 stations is 2 km. A train covers it in 1 minute. It accelerates and retards at the same rate of 0.2m/s². What is the maximum velocity attained by the train?: The distance between 2 stations is 2 km. A train covers it in 1 minute. It accelerates and retards at the same rate of 0.2m/s². What is the maximum velocity attained by the train?
1
Expert's answer
2019-08-12T09:53:41-0400

Lets us find out the equation of maximum velocity when a train first accelerates with "\alpha \space m/sec^2"

and then retards with "\beta \space m/sec^2" with total time being "t" .

for acceleration time "t_1" :

"v=u+at_1"

This v will be "v_{max}=\alpha t_1" .......(1) as u=0

For retardation time t2:

"v=u+at \implies 0=\alpha t_1-\beta t_2" (as final velocity is zero)

"\alpha t_1=\beta t_2
\implies t_2=\frac{ \alpha}{ \beta}\times t_1"

Total time="t=t_1+t_2\implies t_1+\frac{ \alpha}{ \beta}\times t_1=t"

"t_1=\frac{\beta t}{\alpha +\beta}".......(2)

Using (1) and (2)

"v_{max}=\frac{\alpha \beta t}{\alpha +\beta}"

In this question "\alpha ,\beta" and "t" are 0.2"m\/sec^2" ,0.2"m\/sec^2" and 60 sec(1 min=60 sec) respectively.


"v_{max}=\frac{0.2\times0.2\times60}{0.2+0 .2}=6m/sec"